# Adaptive Paradigm Synergy: Can a Cross-Paradigm Objective Enhance Long-Tailed Learning?

### v1

@@ 3.4 Explicit Quantity Awareness
In Sec. 3.3, we describe how we use dynamic temperature tuning strategy to optimize spatial structure. However, we find that in order to increase the stability of training and prevent overfitting, the temperature setting has a certain limit range. At the same time, adjusting the temperature parameter according to different datasets is computationally expensive, so a fixed temperature range is a common choice. This means that when facing real-world datasets with large imbalance factors, our method may not achieve the desired correction effect.
This under-optimization is mainly manifested in the destruction of category-level uniformity due to temperature restrictions. However, this category-level uniformity information can be directly obtained from the sample quantity distribution. Therefore, we introduce a classic supervised re-weighting strategy into self-supervised learning paradigm to assist category homogenization. This supervision can work together with temperature adjustment. At the same time, in order to avoid over-correction, we use square root inverse proportional correction instead of general inverse proportional correction.
3.5 Training with Hybrid Cross-paradigm Loss
@@
$\ell_{i}=-\log\frac{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau_{i})}{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau_{i})+\sum_{v_{j}\in\mathcal{B}^{-}_{i}}\frac{1}{\sqrt{|D_{j}|}}S(\bm{v}_{i},\bm{v}_{j},\tau_{i})}~{}$ (5)
@@
Algorithm 1 Pseudo-code of our proposed APS.
@@
1:  Warm up model $\theta_{0}$ for $B$ epochs according to SimCLR.
@@
6:        Obtain the temperature $\mathbf{t}$ by $P$
7:        Compute the proposed $\mathcal{L}_{\mathrm{APS}}$.
8:        Uptate model $\theta$ by $\mathcal{L}_{\mathrm{APS}}$.
@@ 3.6 More Discussion
Other Self-supervised Framework also could Benefit from our APS. Our APS can be combined with other contrastive learning frameworks [16, 19] or non-contrastive self-supervised frameworks [18], even be transferred into other modalities. Correcting potential imbalances by using only temporary image representations results in high availability and low overhead. The feasibility and extensibility of APS have been validated by combining it with SimCLR, and we will make combining it with other self-supervised learning frameworks as the centre of our future work!
@@
APS 64.17 62.83 74.16 74.27 68.15 85.46
Table 1: Performance comparison on CIFAR10-LT dataset. We utilizes Resnet18 and Resnet50 as backbone to train some recent long-tailed self-supervised methods, and we bold the best performance.
4 Experiment
4.1 Datasets
CIFAR10-LT and CIFAR100-LT[12] are long-tailed versions of the CIFAR10 and CIFAR100 datasets, respectively. The original CIFAR10[27] dataset comprises of 32×32 images from 10 classes, with 50k images allocated for training and 10k images for testing. The imbalance factor is defined as the number of the head classes divided by the tail classes. Following SDCLR[22] and TS[28], we set the imbalance factor to 100. Each class contains images from 4500 to 45 according to the Pareto distribution, a total of 11165 images for training and test set is remained at 10k images.
Similarly, the CIFAR100 dataset has 100 classes of images with 50k images for training and 10k for testing. The number of images per class for training in CIFAR100-LT varies from 450 to 4, with a total of 9,754 images. we average all of our results three times over five splits to reduce randomness to the greatest extent possible..
ImageNet100-LT is a subset of ImageNet100[40] dataset, consisting of 12.21k images from 100 classes for training and 5k images for testing. The images of each class in the training set range from 1280 to 5 according to the Pareto distribution.
4.2 Experimental Setup
Evaluation. The previous exploration, TS[28] provides a comprehensive validation that employs K nearest neighbours[10] (KNN) to verify the latent space structure by directly evaluating the learned local representations, and two linear probing (LP) methods to validate the generalization and linear separability[22] of learned embedding, which we call the balanced minimum sampling linear probing (MS LP) and the long-tailed linear probing (LT LP). We follow their validation methods, and in addition, we provide two other more widely used validation methods, the 1% sampling linear probing (MS LP) and the full set linear probing (Full LP), all of which we will explain in detail below.
@@ 4.2 Experimental Setup
APS 32.14 31.34 29.19 29.17 36.73 59.57
Table 2: Performance comparison on CIFAR100-LT dataset.
As for our unique hyperparameters, number of clusters $K$ is set as equal to the number of classes contained in the dataset, i.e., $K$=10 for CIFAR10-LT and 100 for ImageNet100-LT and CIFAR100-LT. The value of B is equivalent to the epoch when clustering begins, i.e. B=50 for CIFAR10-LT and ImageNet100-LT, 300 for CIFAR100-LT. In all three datasets, we assign the frequent of clustering F to 10. Epochs of warming S=500 for CIFAR10-LT, 1000 for CIFAR100-LT, and 200 for ImageNet100-LT. We will have an in-depth analysis of parameters sensitivity and detailed experimental results in Sec. 4.4.
To gain a more detailed insight of the quantitative impact of our method for calibrating domination of majority classes. We divide all classes of each dataset into three groups from highest to lowest sample size: Head, Mid and Tail. For CIFAR10-LT and CIFAR100-LT datasets, we split the three groups evenly, 4:3:3 for CIFAR10-LT and 34:33:33 for CIFAR100-LT. As for ImageNet100-LT dataset, following [31], defining head classes as classes with over 100 samples, and tail as classes under 20 samples, classes have samples from 20 to 100 are assigned as Mid classes.
4.3 Experimental Results
In this section, we summarize the comprehensive experimental results by comparing our APS with recent methods across a range of benchmarks, and offer insights into the observed outcomes.
In the CIFAR10-LT and CIFAR100-LT datasets, we conduct comparison experiments across all six benchmarks previously mentioned. As shown in Tab. 1 and Tab. 2, for comparison of latent space structure with the ResNet18 backbone model, i.e., KNN@1 and KNN@10, we achieve a significant consistent improvement compared to previous methods, demonstrate that our approach efficiently optimises local embedding representations, which are based on the learned metric. In terms of practical considerations, we adopt three settings of linear probing experiment: 1%S LP, MS LP and LT LP. We find that the average accuracy is positively correlated with the number of fine-tuned samples. For the MS LP and 1%S LP benchmarks, the proposed method continues to demonstrate satisfactory performance when utilising a few-shot setting. The distinction in uniformity yielded by the two settings exerts the most significant influence on SimCLR, while has a limit impact on other methods. At setting of LT LP, we utilize original long-tailed pre-training set for fine-tune. Compare with MS LP and 1%S LP, LT LP has a greater uniformity and more fine-tuning instances. Despite the reduction in the lead, our APS still achieves a competitive result. But for the settings of Full LP, our APS does not perform as expected, especially in the CIFAR100-LT dataset.
In order to gain a deeper understanding of whether this performance discrepancy persists under different alternative models, a series of analogous experiments were conducted around ResNet50. We summarise the experimental results in the ResNet50 section of Tab. 1 and Tab. 2. The experimental results indicate that our methodology does not result in performance collapse when utilising ResNet50, maintains an excellent performance like KNN series in settings of ResNet18 backbone. It is worth noting that both temperature adjustment methods (i.e., TS and APS) yielded the similar performance collapse, and this crash appeared to be exclusive to the CIFAR100-LT dataset, which features a greater number of categories, and was more pronounced when utilising more fine-tuned data.
Method $\bm{\tau}$ $\bm{w}$ KNN@1 KNN@10 1%S LP MS LP LT LP Full LP
[…]
@@ 4.3 Experimental Results
APS-800 39.54 40.26 45.48 47.82 68.30
APS-1700 41.06 42.06 47.66 48.12 69.10
Table 4: Performance comparison on ImageNet100-LT dataset. The best performance is highlighted in bold. APS-800 denotes the result of training our method for 800 epochs of fair comparison. APS-1700 train our method for 1700 epochs to verify scaling performance marked in italics.
Further, we also summary performance comparison on ImageNet100-LT dataset with ResNet50, as shown in Tab. 4. We trained two versions of 800 epochs for fair comparison and 1700 epochs for scale testing. The results show that our APS still achieves superior performance and benefit from increased training time.
@@
4.4 Analysis of sensitivity
We verify the hyperparameters sensitivity of our APS. Firstly, as we simply utilize K-Means[32] for assigning the dynamic labels, the number of clusters $\bm{K}$ should be given carefully consideration, it is challenging to ascertain the total number of image categories without ground truth labels, and finding the optimal parameters is computationally expensive and time-consuming working on large-scale datasets. However, the strong robustness exhibited by proposed APS will alleviate this problem, as shown in Tab. 5. We validate performance of various $\bm{K}$ on CIFAR100-LT dataset, the best item is indicated by bold text and underline for the worst one. Additionally, we report the accuracy range of each benchmark, spanning from 0.27% to 1.23%. Our method maintains competitive results even with K=200, and the high robustness of $K$ helps us to find the optimal parameters more efficiently in the face of unknown datasets.
@@ 4.4 Analysis of sensitivity
Furthermore, we further conduct sensitivity experiment of $\bm{B}$, which determines when to start using the clustering results to influence the learning objective. Setting $B$ larger may harm the concentration on long-tailed learning process, while too small may have impact on the clustering results and drive the model to learn from non-ideal metrics, As shown in Tab. 6. We fixed $K$=10 for experiments on the CIFAR10-LT dataset (split 1). From the results, we can see that the performance difference in $B$ variation from 0 to 500 is minimal with the setup of 2000 epochs training.
$F$ $S$ 200 500 800 1000 $R_{F}$
@@
Table 7: The 1%S LP performance of varying $\bm{F}$ and $\bm{S}$ on CIFAR10-LT dataset with ResNet18. $\bm{F}$ denotes the frequent of clustering, while $\bm{S}$ denotes epochs of warming. The best performance is highlighted in bold. $R_{F}$ and $R_{S}$ denote the range of fluctuations of $F$ and $S$ respectively.
Beyond that, we also conduct more sensitivity validation of clustering hyper-parameters shown as Tab. 7, i.e., clustering frequency $F$ and clustering warming epochs $S$. We report the 1%S LP performance on CIFAR10-LT split 1 with $K$=10 and $B$=50, and bold the best accuracy. The results demonstrated that the performance remained consistent across a range of parameter combinations. The proposed APS displays a great robustness in multiple settings, thereby reducing the additional tuning overhead typically associated with transferring to other public datasets.
@@ 4.5 Ablation study
To verify the importance of each component of proposed APS, we conduct ablation experiments for proposed component, as shown in Tab. 3. Our approach has adaptive temperature awareness and quantity-aware re-weighting, both of which utilise the temporary labels obtained from clustering to correct the neglect of tail classes.
Compared to SimCLR, our method maintains a consist gain from clustering information. APS improves the performances on all the benchmarks ranging from 3.97% to 9.26%, with an average improvement of 6.75%. Even only utilising the re-weighting strategy, APS can also have an average of 3.90% improvement.
In contrast to the TS using temperature cosine variation strategy, APS still improves accuracy by about 0.4% on each benchmark when only using temperature adaptation. These results corroborate the effectiveness of our proposed clustering enhancement approach.
@@ 4.6 T-SNE visualization
T-SNE provide a qualitative perspective into the learned discriminative representation. As shown in Fig. 3, we select samples from all 10 classes of CIFAR10-LT to generate plots. Compared to the SDCLR, Our APS exhibits tighter clusters and more clearly distinguishing boundaries. The results demonstrate that our method is effective in producing high-quality auxiliary supervision when facing long-tailed distributions.
@@ 5 Conclusion
In this paper, we proposed a novel method for self-supervised long-tailed recognition called APS. A cross-paradigm collaborative loss function is devised to facilitate the synergy between adaptive temperature adjustment and re-weighting techniques, thereby mitigating the issue of neglecting tail classes and enabling the acquisition of more generalized embeddings. Our approach demonstrates that the experience of supervised long-tail learning can be migrated to self-supervised long-tail learning with low overhead, in contrast to previous work that proposed non-generic solutions due to the absence of labels. We conducted various benchmarks on multiple datasets to demonstrate the efficiency and robustness of our proposed method. T-SNE[41] visualization analysis also provides a perspective showing better separability of our local representations. In the future, we will explore extending APS to more non-contrast learning SSL frameworks e.g., MAE[18].

### v2

@@ 3.4 Explicit Quantity Awareness
In Sec. 3.3, we describe how we use dynamic temperature tuning strategy to optimize spatial structure. However, we find this to be a flaw in practice, as the general temperature settings are treated as fixed hyperparameters, and pseudo-labels cannot provide information on imbalance ratios. In other words, this flaw gives rise to the fact that the same temperatures will always be assigned to the head class samples on different datasets with different imbalanced ratio. Therefore, we require an adjunct that can take advantage of the information on the distribution of pseudo-labels to jointly serve as a hybrid objective.
Therefore, we introduce a classic supervised re-weighting strategy to assist category uniformization. For each sample, as negative examples, a weight is added inversely proportional to the number of samples in the category to which they belong. And to prevent overcorrection we use the inverse of the square root of the sample size as the weights.
3.5 Training via TASE
@@
$\ell_{i}=-\log\frac{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau_{i})}{S(\bm{v}_{i},\bm{v}_{i^{\prime}},\tau_{i})+\sum_{\bm{v}_{j}\in\mathcal{B}^{-}_{i}}\frac{1}{\sqrt{|D_{j}|}}S(\bm{v}_{i},\bm{v}_{j},\tau_{i})}~{}$ (5)
$\mathcal{L}_{TASE}=\frac{1}{|D|}\sum_{\bm{v}_{i}\in D}{\ell_{i}}$ (6)
@@
Algorithm 1 Pseudo-code of our proposed TASE.
@@
1:  Warm up model $\theta_{0}$ for $B$ epochs via SimCLR.
@@
6:        Obtain the temperature $\mathbf{t}$ for each sample by $P$
7:        Compute the proposed $\mathcal{L}_{\mathrm{TASE}}$.
8:        Uptate model $\theta$ by $\mathcal{L}_{\mathrm{TASE}}$.
@@ 3.6 More Discussion
Expandability Our TASE can be combined with other contrastive learning frameworks [15, 17] or non-contrastive self-supervised frameworks [18]. Correcting potential imbalances by using only temporary image representations results in high availability and low overhead. The feasibility and extensibility of TASE have been validated by combining it with SimCLR, and we will make combining it with other self-supervised learning frameworks in the future.
4 Experiment
4.1 Datasets
CIFAR10-LT and CIFAR100-LT [11] are long-tailed versions of the CIFAR10 and CIFAR100 datasets, respectively. The original CIFAR10 [25] dataset comprises of 32×32 images from 10 classes, with 50k images allocated for training and 10k images for testing. The imbalance factor is defined as the number of the head classes divided by the number of tail classes. Following SDCLR [20] and TS [26], we set the imbalance factor to 100. Each class contains between 4,500 and 45 images according to the Pareto distribution. In total, there are 11,165 images in the training and test sets, with 10,000 images remaining.
Similarly, the CIFAR100 dataset has 100 classes of images with 50k images for training and 10k for testing. The number of images per class for training in CIFAR100-LT varies from 450 to 4, with a total of 9,754 images. We average all of our results over three runs across five splits to reduce randomness to the greatest extent possible.
ImageNet100-LT is a subset of ImageNet100 [37] dataset, consisting of 12.21k images from 100 classes for training and 5k images for testing. The number of images of each class in the training set range from 1280 to 5 according to the Pareto distribution.
4.2 Experimental Setup
Evaluation. TS [26] provides a comprehensive validation that employs K nearest neighbours [9] (KNN) to verify the latent space structure by directly evaluating the learned local representations, and two linear probing (LP) methods to validate the generalization and linear separability [20] of learned embedding, which we call the balanced minimum sampling linear probing (MS LP) and the long-tailed linear probing (LT LP). We follow their validation methods, and in addition, we provide two other more widely used validation methods, the 1% sampling linear probing (1%S LP) and the full set linear probing (Full LP), all of which we will explain in detail below.
@@ 4.2 Experimental Setup
TASE 64.17 62.83 74.16 74.27 68.15 85.46
Table 1: Performance comparison on CIFAR10-LT dataset. We utilizes Resnet18 and Resnet50 as backbone to reproduce some recent long-tailed self-supervised methods, and we bold the best performance.
As for our unique hyperparameters, The value of clusters $K$ is set as equal to the number of classes contained in the dataset, i.e., $K$=10 for CIFAR10-LT and 100 for ImageNet100-LT and CIFAR100-LT. The value of B is equivalent to the starting epoch of clustering, i.e. B=50 for CIFAR10-LT and ImageNet100-LT, 300 for CIFAR100-LT. In all three datasets, we assign the frequent of clustering F to 10. Epochs of warming S=500 for CIFAR10-LT, 1000 for CIFAR100-LT, and 200 for ImageNet100-LT. We will have an in-depth analysis of parameters sensitivity and detailed experimental results in Sec. 4.4.
To gain a more detailed insight of the quantitative impact of our method for calibrating domination of majority classes, we divide all classes of each dataset into three groups from highest to lowest sample size: Head, Mid and Tail. For CIFAR10-LT and CIFAR100-LT datasets, we split the three groups evenly, 4:3:3 for CIFAR10-LT and 34:33:33 for CIFAR100-LT. As for ImageNet100-LT dataset, following [29], defining head classes as classes with over 100 samples, and tail as classes under 20 samples, classes have the number of samples from 20 to 100 are assigned as Mid classes.
4.3 Experimental Results
In this section, we summarize the comprehensive experimental results by comparing our TASE with recent methods across a range of benchmarks, and offer insights into the observed outcomes.
In the CIFAR10-LT and CIFAR100-LT datasets, we conduct comparison experiments across all six benchmarks previously mentioned. As shown in Tab. 1 and Tab. 2, for comparison of latent space structure with the ResNet18 backbone model, i.e., KNN@1 and KNN@10, we achieve a significant consistent improvement compared to previous methods, which demonstrates that our approach efficiently optimises local embedding representations, which are based on the learned metric. In terms of practical considerations, we adopt three settings of linear probing experiment: 1%S LP, MS LP and LT LP. We find that the average accuracy is positively correlated with the number of fine-tuned samples. For the MS LP and 1%S LP benchmarks, the proposed method continues to demonstrate satisfactory performance. Differences in data intersections for training and fine-tuning yielded by the two settings exerts the most significant influence on SimCLR, while has a limit impact on other methods. At setting of LT LP, we utilize original long-tailed pre-training set for fine-tune. Compared to MS LP and 1%S LP, LT LP has a greater uniformity and more fine-tuning instances. Despite the reduction in the lead, our TASE still achieves a competitive result. But for the settings of Full LP, our TASE does not perform as expected, especially in the CIFAR100-LT dataset.
@@ 4.3 Experimental Results
TASE 32.14 31.34 29.19 29.17 36.73 59.57
Table 2: Performance comparison on CIFAR100-LT dataset. We utilizes Resnet18 and Resnet50 as backbone to reproduce some recent long-tailed self-supervised methods, and we bold the best performance.
@@
TASE 39.54 40.26 45.48 47.82 68.30
Table 3: Performance comparison on ImageNet100-LT dataset. The best performance is highlighted in bold.
In order to gain a deeper understanding of whether this performance discrepancy persists under different alternative models, a series of analogous experiments were conducted around ResNet50. We summarise the experimental results in the ResNet50 section of Tab. 1 and Tab. 2. The experimental results indicate that our method does not result in performance collapse when utilising ResNet50, and maintains an excellent performance like KNN series in settings of ResNet18 backbone. It is worth noting that both temperature adjustment methods (i.e., TS and TASE) yield the similar performance collapse, and this crash appears to be exclusive to the CIFAR100-LT dataset, which features a greater number of categories, and was more pronounced when utilising more fine-tuned data.
@@
Further, we also summary performance comparison on ImageNet100-LT dataset with ResNet50, as shown in Tab. 3.
@@
4.4 Analysis of Sensitivity
We verify the hyperparameters sensitivity of our TASE. Firstly, as we simply utilize K-means [30] for assigning the dynamic labels, the number of clusters $\bm{K}$ should be given carefully consideration. It is challenging to ascertain the total number of image categories without ground truth labels, and finding the optimal parameters is computationally expensive and time-consuming working on large-scale datasets. However, the strong robustness exhibited by proposed TASE will alleviate this problem, as shown in Tab. 4. We validate performance of various $\bm{K}$ on CIFAR100-LT dataset, the best item is indicated by bold text and underline for the worst one. Additionally, we report the accuracy range of each benchmark, spanning from 0.27% to 1.23%. Our method maintains competitive results even with K=200, and the high robustness of $K$ helps us to find the optimal parameters more efficiently in the face of unknown datasets.
Furthermore, we further conduct sensitivity experiment of $\bm{B}$, which determines when to start using the clustering results to influence the learning objective. Setting $B$ larger may harm the concentration on long-tailed learning process, while too small may have impact on the clustering results and drive the model to learn from non-ideal metrics, as shown in Tab. 5. We fix $K$=10 for experiments on the CIFAR10-LT dataset (split 1). From the results, we can see that the performance difference in $B$ variation from 0 to 500 is minimal with the setup of 2000 epochs training.
Method $\bm{\tau}$ $\bm{w}$ KNN@1 KNN@10 1%S LP MS LP LT LP Full LP
[…]
$\bm{F}$ $\bm{S}$ 200 500 800 1000 $R_{F}$
@@ 4.4 Analysis of Sensitivity
Table 7: The 1%S LP performance of varying $\bm{F}$ and $\bm{S}$ on CIFAR10-LT dataset with ResNet18. $\bm{F}$ denotes the frequent of clustering, while $\bm{S}$ denotes epochs of warming. The best performance is highlighted in bold. $R_{F}$ and $R_{S}$ denote the range of fluctuations of accuracy with fixed $\bm{F}$ and $\bm{S}$, respectively.
Beyond that, we also conduct more sensitivity validation of clustering hyper-parameters shown as Tab. 7, i.e., clustering frequency $F$ and clustering warming epochs $S$. We report the 1%S LP performance on CIFAR10-LT split 1 with $K$=10 and $B$=50, and bold the best accuracy. The results demonstrate that the performance remained consistent across a range of parameter combinations. The proposed TASE displays a great robustness across multiple settings, thereby reducing the additional tuning overhead typically associated with transferring to other public datasets.
@@ 4.5 Ablation Study
To verify the importance of each component of proposed TASE, we conduct ablation experiments for proposed component, as shown in Tab. 6. Our approach has adaptive temperature awareness and quantity-aware re-weighting, both of which utilise the temporary labels obtained from clustering to correct the neglect of tail classes.
Compared to SimCLR, our method maintains a consist gain from clustering information. TASE improves the performances on all the benchmarks ranging from 3.97% to 9.26%, with an average improvement of 6.75%. Even only utilising the re-weighting strategy, TASE can also have an average of 3.90% improvement.
In contrast to the TS using temperature cosine variation strategy, TASE still improves accuracy by about 0.4% on each benchmark when only using temperature adaptation. These results corroborate the effectiveness of our proposed clustering enhancement approach.
@@ 4.6 T-SNE Visualization
T-SNE provides a qualitative perspective into the learned discriminative representation. As shown in Fig. 3, we select samples from all 10 classes of CIFAR10-LT to generate plots. Compared to the SDCLR, our TASE exhibits tighter clusters and more clearly distinguishing boundaries. The results demonstrate that our method is effective in producing high-quality auxiliary supervision when facing long-tailed distributions.
@@ 5 Conclusion
In this paper, we proposed a novel method for self-supervised long-tailed recognition called TASE. We introduce pseudo-label assignment to drive the dynamic temperature strategy and re-weighting to counteract the long-tailed distribution, training different samples with optimal training modes to obtain robust representations. Our approach demonstrates that the experience of supervised long-tail learning can be migrated to self-supervised long-tail learning with low overhead, in contrast to previous work that proposed non-generic solutions due to the absence of labels. We conduct various benchmarks on multiple datasets to demonstrate the efficiency and robustness of our proposed method. T-SNE [39] visualization analysis also provides a perspective showing better separability of our local representations.
Supplementary Material
6 Other Experimental Results
6.1 Fine-grained Results
We also summarize a detailed results of CIFAR10-LT and ImageNet100-LT to verify the performance of fine-grained classes in Tab. 10 and Tab. 11.
We provide accuracy for different categories.The situation of split is referenced in Sec. 4.2.
6.2 Results on Balanced Datasets
We also performed the same experiments on the balanced dataset, as shown in Tab. 8.
The results show that our method can achieve a significant lead on most benchmarks, and a similar performance to the previous work on Full LP. This indicates that our method is effective in identifying potential subclasses within a balanced dataset, thereby enhancing the performance.
Method K@1 K@10 MS LP 1%S LP LT LP Full LP
SimCLR 53.14 47.15 67.86 74.81 61.85 80.69
SDCLR 64.01 60.50 76.38 75.81 71.42 84.75
TS 65.33 62.95 76.99 78.56 69.70 83.32
TASE 69.61 67.44 78.27 79.02 71.77 84.15
Table 8: Performance comparison on balanced CIFAR10 dataset with ResNet18.
6.3 Results via Other Clustering Methods
Alternative clustering methods are employed to implement our TASE, result is summarized in Tab. 9. Experimental results indicate that K-means and agglomerative clustering exhibit comparable performance. More clustering methods such as discriminative clustering will be explored in future work.
Method K@1 K@10 MS LP 1%S LP LT LP Full LP
DBSCAN 60.96 60.77 69.74 68.97 61.83 78.67
A. C. 62.22 61.49 71.57 71.60 64.39 79.46
K-means 63.08 62.30 71.78 70.84 64.50 79.58
Table 9: Effect of different clustering methods on CIFAR10-LT dataset (Split 1) with ResNet18. A. C. denotes Agglomerative Clustering.
7 Datasets Source
We used three datasets: CIFAR10-LT, CIFAR100-LT, and ImageNet100-LT. They were all obtained by long-tail subset sampling from the original datasets, namely CIFAR10, CIFAR100, and ImageNet. Sources of original datasets can be found in Tab. 12. These datasets are publicly available and widely used. All data are in full compliance with all applicable ethical standards.
8 Computing Infrastructure
Our codes are based on Pytorch 2.4.0. Our models and reproductions of previous work are trained on GeForce RTX 3090 and GeForce RTX 4090.
KNN@1 KNN@10
Method Head Mid Tail All Head Mid Tail All
SimCLR 87.68$\pm$4.28 62.46$\pm$12.35 29.01$\pm$15.3 62.51 90.68$\pm$4.09 62.14$\pm$12.77 21.91$\pm$18.21 61.49
SDCLR 86.92$\pm$4.25 62.36$\pm$12.15 30.38$\pm$13.37 62.59 90.71$\pm$4.24 61.50$\pm$12.49 21.78$\pm$16.83 61.27
TS 88.39$\pm$3.99 61.91$\pm$13.02 30.59$\pm$14.46 63.11 91.46$\pm$3.84 61.26$\pm$12.60 23.26$\pm$17.07 61.94
TASE 88.63$\pm$3.85 64.15$\pm$11.74 31.58$\pm$14.09 64.17 91.86$\pm$3.66 63.11$\pm$12.03 23.84$\pm$17.50 62.83
MS LP 1%S LP
Method Head Mid Tail All Head Mid Tail All
SimCLR 79.48$\pm$5.30 70.49$\pm$11.73 66.32$\pm$12.48 72.83 79.63$\pm$4.12 71.31$\pm$14.46 63.09$\pm$11.5 72.17
SDCLR 77.75$\pm$6.21 73.10$\pm$10.82 67.57$\pm$13.71 73.30 77.11$\pm$5.97 71.91$\pm$13.98 67.71$\pm$11.34 72.73
TS 79.15$\pm$9.23 73.65$\pm$12.08 65.97$\pm$15.20 73.54 79.38$\pm$5.72 72.51$\pm$13.32 65.93$\pm$11.21 73.28
TASE 79.61$\pm$7.34 73.27$\pm$11.52 67.79$\pm$10.43 74.16 80.30$\pm$5.23 73.53$\pm$12.57 66.96$\pm$12.41 74.27
LT LP Full LP
Method Head Mid Tail All Head Mid Tail All
SimCLR 93.04$\pm$2.95 66.89$\pm$11.95 28.34$\pm$14.20 65.78 87.60$\pm$4.67 83.35$\pm$6.92 79.01$\pm$7.77 83.75
SDCLR 93.68$\pm$3.24 69.53$\pm$9.80 29.15$\pm$13.37 67.07 89.50$\pm$3.42 85.20$\pm$6.00 80.77$\pm$7.01 85.59
TS 93.70$\pm$2.73 68.98$\pm$11.82 28.75$\pm$12.30 66.80 88.00$\pm$4.54 84.41$\pm$6.65 80.79$\pm$6.31 84.76
TASE 94.14$\pm$2.52 70.77$\pm$9.30 30.86$\pm$12.94 68.15 89.03$\pm$4.23 85.03$\pm$6.69 81.12$\pm$7.05 85.46
Table 10: Detailed results on CIFAR10-LT with ResNet50.
KNN@1 KNN@10 1%S LP
Method Head Mid Tail All Head Mid Tail All Head Mid Tail All
SimCLR 55.13 30.00 10.71 37.10 58.51 29.70 8.71 38.00 51.79 36.77 30.29 42.16
SDCLR 55.74 27.67 10.15 36.06 60.68 26.00 5.23 37.36 48.31 39.17 36.46 42.38
TS 57.23 30.26 13.14 38.36 60.41 29.53 10.14 38.86 60.41 40.38 33.57 45.18
TASE-800 59.26 31.84 10.92 39.54 64.11 30.24 8.31 40.26 54.16 40.78 37.85 45.48
TASE-1700 61.53 32.90 12.00 41.06 65.63 32.69 8.46 42.06 56.16 43.43 38.77 47.66
LT LP Full LP
Method Head        Mid        Tail         All Head        Mid         Tail         All
SimCLR 67.59       36.47       9.43       44.82 69.54       63.71       59.69       65.46
SDCLR 71.74       36.86       8.00       46.40 70.10       65.04       60.92       66.48
TS 70.67       38.85      10.29      47.26 72.53       56.27       63.69       67.82
TASE-800 71.16       39.88       9.54       47.82 72.74       66.41       62.46       68.30
TASE-1700 71.47       40.41       8.92       48.12 73.79       66.53       65.08       69.10
Table 11: Detailed results on ImageNet100-LT with ResNet50.TASE-800 denotes the result of training our method for 800 epochs of fair comparison. TASE-1700 train our method for 1700 epochs to verify scaling performance marked in italics.
Dataset Source Link
CIFAR10 https://www.cs.toronto.edu/ kriz/cifar-10-python.tar.gz
CIFAR100 https://www.cs.toronto.edu/ kriz/cifar-100-python.tar.gz
ImageNet http://image-net.org/download
Table 12: The source of the dataset we use.